# Each to Their Own: Exploring the Optimal Embedding in RAG

### v1

@@ 2.1 Retrieval Augmented Generation (RAG)
[…]
@@ 2.2 Confidence of LLMs
where $\text{f}()$ is an LLM, $x$ is the prompt, $y_{<i}$ represents the first generated $i-1$ tokens, and $y_{i}$ denotes the $i^{th}$ token. The output of $\text{f}()$ is the probability distribution over each word in the vocabulary, from which $y_{i}$ is sampled according to this probability vector. This process can be viewed as a Markov Decision Process (MDP), where the state consists of the prompt $x$ along with the currently generated tokens $y_{<i}$, and the action space corresponds to the vocabulary, allowing each token to be chosen with a probability given by
@@ 3 Methodology
Based on the understanding that different embedding models excel in their respective domains, our study aims to develop a comprehensive RAG method that combines the strengths of various embedding models. In this work, given a question $q$, we utilize a single LLM $\text{f}()$ for answer generation, while $N$ embedding models $\{\text{g}_{1}(),\text{g}_{2}(),\ldots,\text{g}_{N}()\}$ assist in retrieving question-related information from a corpus, which has been divided into chunks $C=[c_{1},c_{2},\ldots,c_{m}]$. Inspired by the stacking and bagging paradigms of ensemble learning, we propose two distinct methods: (1) Mixture-Embedding RAG, which first combines the retrieved results from multiple embedding models and incorporates them into the prompt for the LLM, and (2) Confident RAG, which employs vanilla RAG with different embedding models for initial answer generation and subsequently selects the answer with the highest confidence as the final response.
@@ 3.2 Confident RAG
Inspired by Kang et al. (2025), we first let the LLM f generate $N$ separate answers, using the retrieved chunks from the $N$ embedding models separately. Then, we evaluate the confidence of the $N$ answers using the metrics mentioned in Section 2.2. Finally, we select the answer with the highest confidence as our final answer for each question.
@@ 4.2.1 Vanilla LLM and RAG
LLM Model Vanilla LLM Vanilla RAG Emb1 Emb2 Emb3 Emb4 Avg Improvement Qwen2.5-Math-7B Yang et al. (2024) 75.2% 81.0% 83.0% 77.6% 80.2% 80.5% 5.3$\uparrow$ Llama-3.1-8B Grattafiori et al. (2024) 16.6% 21.8% 19.8% 19.4% 24.0% 21.3% 4.7$\uparrow$ OLMo-2-1124-7B OLMo et al. (2024) 21.0% 26.0% 28.4% 25.0% 24.6% 26.0% 5.0$\uparrow$ Average 37.6% 42.9% 43.7% 40.7% 42.9% 42.6% 5.0$\uparrow$
@@ 4.2.2 Results of Mixture-Embedding RAG method
LLM Model Mix-Embedding RAG 2 Embs 3 Embs 4 Embs Avg v.s. Vanilla LLM v.s. Vanilla RAG Qwen2.5-Math-7B Yang et al. (2024) 74.20% 76.00% 74.80% 75.0% -0.2$\downarrow$ -5.5$\downarrow$ Llama-3.1-8B Grattafiori et al. (2024) 20.90% 19.60% 22.80% 21.1% 4.5$\uparrow$ -0.2$\downarrow$ OLMo-2-1124-7B OLMo et al. (2024) 26.20% 26.80% 26.40% 26.5% 5.5$\uparrow$ 0.5$\uparrow$ Average 40.4% 40.8% 41.3% 40.9% 3.3$\uparrow$ -1.7$\downarrow$
@@ 4.2.3 Results of Confident RAG method
LLM Emb. Model AvgLogP Self-certainty Gini Entropy DP Qwen2.5-Math-7B Yang et al. (2024) 1,2 82.0% 85.0% 81.8% 82.8% 83.6% 1,3 79.4% 83.4% 79.2% 79.6% 81.0% 1,4 81.4% 84.6% 81.6% 81.8% 82.2% 2,3 81.8% 84.4% 81.4% 81.8% 82.4% 2,4 81.4% 83.6% 81.0% 81.4% 82.2% 3,4 79.2% 82.2% 79.8% 79.8% 80.6% Avg (n=2) 80.9% 83.9% 80.8% 81.2% 82.0% 1,2,3 79.4% 85.0% 79.0% 79.8% 81.6% 1,2,4 79.6% 85.0% 79.6% 80.6% 82.0% 1,3,4 79.0% 84.8% 79.6% 80.0% 80.8% 2,3,4 79.2% 84.4% 79.0% 79.8% 81.0% Avg (n=3) 79.3% 84.8% 79.3% 80.1% 81.4% 1,2,3,4 78.2% 84.8% 78.4% 79.2% 80.6% Avg (n=2,3,4) 80.1% 84.3% 80.0% 80.6% 81.6% v.s. Vanilla RAG 0.4$\downarrow$ 3.8$\uparrow$ 0.4$\downarrow$ 0.1$\uparrow$ 1.2$\uparrow$ v.s. Vanilla LLM 4.9$\uparrow$ 9.1$\uparrow$ 4.8$\uparrow$ 5.4$\uparrow$ 6.4$\uparrow$ Llama-3.1-8B Grattafiori et al. (2024) 1,2 27.2% 27.2% 27.0% 27.0% 27.2% 1,3 26.8% 26.8% 26.6% 27.0% 27.2% 1,4 26.8% 26.4% 26.4% 26.4% 26.4% 2,3 23.2% 24.0% 24.0% 24.0% 23.6% 2,4 26.6% 26.8% 26.4% 26.0% 26.6% 3,4 25.8% 26.4% 26.6% 26.6% 26.4% Avg (n=2) 26.1% 26.3% 26.2% 26.2% 26.2% 1,2,3 28.6% 29.2% 28.8% 29.4% 29.4% 1,2,4 28.8% 28.6% 28.0% 28.0% 28.2% 1,3,4 27.4% 27.4% 27.2% 27.6% 27.6% 2,3,4 25.8% 26.4% 26.0% 26.0% 26.4% Avg (n=3) 27.7% 27.9% 27.5% 27.8% 27.9% 1,2,3,4 27.8% 27.6% 27.0% 27.4% 27.6% Avg (n=2,3,4) 26.8% 27.0% 26.7% 26.9% 27.0% v.s. Vanilla RAG 5.6$\uparrow$ 5.7$\uparrow$ 5.5$\uparrow$ 5.6$\uparrow$ 5.7$\uparrow$ v.s. Vanilla LLM 10.2$\uparrow$ 10.4$\uparrow$ 10.1$\uparrow$ 10.3$\uparrow$ 10.4$\uparrow$ OLMo-2-1124-7B OLMo et al. (2024) 1,2 31.0% 31.2% 30.4% 31.2% 32.2% 1,3 29.8% 29.6% 29.8% 30.0% 30.0% 1,4 29.4% 29.2% 28.6% 29.6% 31.0% 2,3 31.6% 30.8% 30.2% 31.4% 32.8% 2,4 32.6% 32.0% 31.0% 33.0% 33.8% 3,4 29.6% 30.2% 29.6% 30.6% 31.4% Avg (n=2) 30.7% 30.5% 29.9% 31.0% 31.9% 1,2,3 32.6% 32.0% 31.4% 32.2% 33.2% 1,2,4 32.8% 32.8% 31.6% 33.2% 34.8% 1,3,4 31.2% 31.2% 30.6% 32.0% 34.6% 2,3,4 32.6% 32.2% 30.8% 33.6% 36.6% Avg (n=3) 32.3% 32.1% 31.1% 32.8% 34.8% 1,2,3,4 32.8% 32.4% 31.0% 33.2% 35.8% Avg (n=2,3,4) 31.5% 31.2% 30.5% 31.8% 33.3% v.s. Vanilla RAG 5.5$\uparrow$ 5.2$\uparrow$ 4.5$\uparrow$ 5.8$\uparrow$ 7.3$\uparrow$ v.s. Vanilla LLM 10.5$\uparrow$ 10.2$\uparrow$ 9.5$\uparrow$ 10.8$\uparrow$ 12.3$\uparrow$ Average Avg (n=2) 45.9% 46.9% 45.6% 46.1% 46.7% Avg (n=3) 46.4% 48.3% 46.0% 46.9% 48.0% Avg (n=4) 46.3% 48.3% 45.5% 46.6% 48.0% Avg (n=2,3,4) 46.1% 47.5% 45.7% 46.4% 47.3% v.s. Vanilla RAG 3.5$\uparrow$ 4.9$\uparrow$ 3.2$\uparrow$ 3.9$\uparrow$ 4.7$\uparrow$ v.s. Vanilla LLM 8.5$\uparrow$ 9.9$\uparrow$ 8.1$\uparrow$ 8.8$\uparrow$ 9.7$\uparrow$

### v2

@@ 2.1 Retrieval Augmented Generation (RAG)
where $s_{q,c_{j}}$ represents the cosine similarity between question $q$ and chunk $c_{j}$, $\cdot$ denotes the dot product, and $\|\cdot\|$ denotes the Euclidean norm. Finally, we select $k$ chunks with the highest similarity and incorporate them into the prompt, expressed as:
@@ 2.2 Confidence of LLMs
where $\text{f}()$ is an LLM, $x$ is the prompt, $y_{<i}$ represents the first generated i1i-1 tokens, and $y_{i}$ denotes the $i^{th}$ token. The output of $\text{f}()$ is the probability distribution over each word in the vocabulary, from which $y_{i}$ is sampled according to this probability vector. This process can be viewed as a Markov Decision Process (MDP), where the state consists of the prompt $x$ along with the currently generated tokens $y_{<i}$, and the action space corresponds to the vocabulary, allowing each token to be chosen with a probability given by
@@ 3 Methodology
Based on the understanding that different embedding models excel in their respective domains, our study aims to develop a comprehensive RAG method that combines the strengths of various embedding models. In this work, given a question $q$, we utilize a single LLM $\text{f}()$ for answer generation, while NN embedding models $\{\text{g}_{1}(),\text{g}_{2}(),\ldots,\text{g}_{N}()\}$ assist in retrieving question-related information from a corpus, which has been divided into chunks $C=[c_{1},c_{2},\ldots,c_{m}]$. Inspired by the stacking and bagging paradigms of ensemble learning, we propose two distinct methods: (1) Mixture-Embedding RAG, which first combines the retrieved results from multiple embedding models and incorporates them into the prompt for the LLM, and (2) Confident RAG, which employs vanilla RAG with different embedding models for initial answer generation and subsequently selects the answer with the highest confidence as the final response.
@@ 3.2 Confident RAG
Inspired by Kang et al. (2025), we first let the LLM f generate NN separate answers, using the retrieved chunks from the NN embedding models separately. Then, we evaluate the confidence of the NN answers using the metrics mentioned in Section $2.$ Finally, we select the answer with the highest confidence as our final answer for each question.
@@ 4.2.1 Vanilla LLM and RAG
[…]
@@ 4.2.2 Results of Mixture-Embedding RAG method
[…]
@@ 4.2.3 Results of Confident RAG method
[…]